OpenStudy (anonymous):

Let f(x,y,z)=x+ln(2y+3z^2) Evaluate ∂^2f/∂y∂x (2,1,-2)

OpenStudy (anonymous):

∂^2f/∂y∂x=(∂f/∂y)x(∂f/∂x) ∂f/∂y=2/(2y+3z^2) ∂f/∂x=1 so ∂^2f/∂y∂x=2/(2y+3z^2)=1/5

OpenStudy (anonymous):

thanks i thought it was 0

OpenStudy (anonymous):

your welcme

OpenStudy (anonymous):

is it 1/7??